OpenStudy (anonymous):

Letf(x)= 9x2+4 . (2x + 1)(x − 2)2 (i) Express f(x) in partial fractions. (ii) Show that, when x is sufficiently small for x3 and higher powers to be neglected, f(x) = 1 − x + 5x2. What does it mean to show that x is sufficiently small for x^3 ?

OpenStudy (anonymous):

is this supposed to be a fraction?

OpenStudy (anonymous):

The first part yes, partial fraction, but I have the answers You want them?

OpenStudy (anonymous):

i don't understand the expression can you write it using the equation editor?

OpenStudy (anonymous):

@satellite73 Let \[f(x)=\frac{ 9x2+4 }{ (2x+1)(x−2)^2 }\] i) Express f(x) in partial fractions. (ii) Show that, when x is sufficiently small for x3 and higher powers to be neglected, \[f(x)=1−x+5x^2.\]

OpenStudy (anonymous):

lord is this ugly

OpenStudy (anonymous):

i got the answer using wolfram it is just a pain to do by hand www.wolframalpha.com/input/?i=(9x^2%2B4)%2F((2x-1)(x-2)^2

OpenStudy (anonymous):

I have the answer for a.

OpenStudy (anonymous):

I just need to understand what they want me to do for b

OpenStudy (anonymous):

Should I give you the anwers for a?

OpenStudy (anonymous):

@satellite73 Answer for a. \[\frac{ A }{ (2x+1) }+\frac{ B }{ (x – 2) }+\frac{ C }{ (x – 2)^2 }\] Where A=1 B=4 C=8 such that \[\frac{ 1 }{ (2x+1) }+\frac{ 4 }{ (x – 2) }+\frac{ 8 }{ (x – 2)^2 }\]

OpenStudy (anonymous):

hold on that is not what i got

OpenStudy (anonymous):

ooh you are right for sure sorry

OpenStudy (anonymous):

for your second question i am at a loss if you ignore the cube term in the denominator you get \[\frac{9x^2+4}{-7x^2+4x+4}\]

OpenStudy (anonymous):

maybe you are supposed to divide backwards

OpenStudy (anonymous):

oh yeah, that works

OpenStudy (anonymous):

use long division \(4+4x-7x^2|4+0x+9x^2\)

OpenStudy (anonymous):

it is very hard for me to write long division here, but that i think will give you what you want set up the long division working from the bottom up instead of from the top down

OpenStudy (anonymous):

But what does it mean to show that x is sufficiently small ?

OpenStudy (anonymous):

it means "if x is sufficiently small ..." not "show that x is sufficiently small"

OpenStudy (anonymous):

Don't worry I get it you expand binomially thanks for your help anyway